Read and analyze the given problem. Provide a complete solution, Table, and a graph corresponds to the answer.


Due to ASF issue it was aired that the average price of a kilo of pork in MM is P195.00. However, a sample of 15 prices randomly collected from different markets in MM showed an average of P200.00 and with standard deviation of P9.50. using 0.05 significant level. What would be your conclusion for the price of pork in MM.














1
Expert's answer
2021-06-07T13:29:02-0400

"\mu=195 \\

n=15 \\

\bar{x}=200 \\

\sigma=9.50 \\

α=0.05 \\

H_0: \mu=195 \\

H_1: \mu>195 \\

The \;test\; statistic: \\

z=\frac{\bar{x}-\mu}{\sigma / \sqrt{n}} \\

= \frac{200-195}{9.50 / \sqrt{15}} \\

= \frac{5}{2.453} \\

=2.03 \\

Z_{crit}=1.645 \\

Z=2.03>Z_{crit}=1.645"

So, we reject H0 and accept H1.

There is evidence to support the claim that the average price of a kilo of pork in MM is more than P195.00.